OpenStudy (anonymous):

140, 142, 142, x, 170, 172 a) If the median of this data set is 147, what is the value of x?

OpenStudy (anonymous):

I'll try

OpenStudy (anonymous):

okidoke!

OpenStudy (anonymous):

I think 116 I have to check not sure

OpenStudy (anonymous):

How did you get that ? :3

OpenStudy (mrnood):

in an even number of samples the median is generally taken to be the mean of the 2 middle numbers So in this case 142 and x are the miidle numbers so the mean of 142 and x is 147. this means (142+x)/2 = 147 You can do it from there....

OpenStudy (anonymous):

Yes it is x=116

OpenStudy (mrnood):

@ttb123456789 that is not correct

OpenStudy (anonymous):

Oh sorry. I conted mean not median Let me redo

OpenStudy (mrnood):

not even the mean is 116

OpenStudy (anonymous):

Why

OpenStudy (anonymous):

I thought you can set it up as (142+x)/2=147

OpenStudy (anonymous):

Wait, MrNood, after 142+x/2 = 147, you multiply both sides by 2 and go through that process right?

OpenStudy (mrnood):

@unicornsandmarshmellows yes - but beware - where I put the brackets

OpenStudy (anonymous):

yeah i notified that thanks !

OpenStudy (anonymous):

alrighty awesome thanks for your help!

OpenStudy (mrnood):

@ttb123456789 I apologise : IF the mean were 147 then x would be 116